OpenStudy (hzro):

Use operations of decimal, fraction, and percent numbers to model the following scenarios. In your final answer, include only an equation or proportion modeling the scenario. Do not solve the problems. 1. A family dines in a popular franchise restaurant. At the end of the meal, they decide to leave their server a monetary tip that is equal to 20% of the total bill amount, $60.50. How much will the family leave their server as a tip?

OpenStudy (welshfella):

20% = 20/100 = 1/5 = 0.20

OpenStudy (hzro):

Thanks can you do another for me? 2. A family dines in a popular franchise restaurant. They plan to use a coupon that will give them a discount of 15% off of their total dinner bill, not including applicable sales tax or their server’s tip. If the family’s dinner bill totaled $34.00 before tax and tip, how much money will they save by using the coupon?